# Efficient and Workload-Aware LLM Serving via Runtime Layer Swapping and KV Cache Resizing

### v1

@@ Abstract
Efficiently serving large language models (LLMs) under dynamic and bursty workloads remains a key challenge for real-world deployment. Existing serving frameworks and static model compression techniques fail to adapt to workload fluctuations, leading to either service-level objective (SLO) violations under full-precision serving or persistent accuracy degradation with static quantization. We present MorphServe, a dynamic, workload-aware LLM serving framework based on morphological adaptation. MorphServe introduces two asynchronous, token-level runtime mechanisms: quantized layer swapping, which selectively replaces less impactful layers with quantized alternatives during high-load periods, and pressure-aware KV cache resizing, which dynamically adjusts KV cache capacity in response to memory pressure. These mechanisms enable state-preserving transitions with minimum runtime overhead and are fully compatible with modern scheduling and attention techniques. Extensive experiments on Vicuna and Llama family models with real-world workloads demonstrate that MorphServe reduces average SLO violations by 92.45% and improves the P95 TTFT latency by $\times2$3.9×3.9$\times$3.9 × compared to full-precision serving, without compromising generation quality. These results establish MorphServe as a practical and elastic solution for LLM deployment in dynamic environments.
@@ 1 Introduction
The rise of large language models (LLMs) has made efficient and reliable serving a core challenge in modern AI infrastructure. Systems like vLLM [33] and TGI [29] optimize throughput via PagedAttention [33] and continuous batching [75, 62, 24], but assume fixed-precision execution and stable workloads. In contrast, real-world LLM workloads are dynamic and bursty [67, 4], with fluctuating request rates and context lengths. Even brief load spikes can cause memory exhaustion or queueing delays, leading to SLO violations—e.g., higher time-to-first-token (TTFT) and time-per-output-token (TOPT)—that degrade user experience and system throughput.
@@
One naive solution is to statically over-provision GPU resources to accommodate worst-case traffic spikes. However, over-provisioning leads to substantial cost inefficiencies during underutilized periods [30, 19]. Moreover, edge deployments lack the flexibility for dynamic scaling altogether [6]. Thus, the inability to elastically match model resource usage to real-time demand results in either SLO violations under pressure, or significant resource waste during low-load intervals.
@@
Model compression techniques, such as quantization [38, 17, 37, 57], pruning [43, 59, 21], or low-rank approximation [26, 70], offer an alternative approach by statically reducing the resource footprint of deployed LLMs. While these methods are effective in lowering memory and compute demands, they introduce irreversible accuracy degradation that persists even during periods of low load, when full-precision inference could be served without penalty. This results in a rigid, suboptimal quality–efficiency tradeoff that fails to align with workload variability. Key-value cache (KVC) compression [7, 78, 35] and eviction [41, 16] methods have been proposed to further reduce memory usage. However, these techniques often rely on fixed heuristics, cannot adapt to different workloads, lack compatibility with modern attention variants like Grouped Query Attention (GQA) [3, 9, 10] and Multi-Head Latent Attention (MLA) [45, 40], and remain inflexible to runtime serving conditions.
@@
In this paper, we present MorphServe, a dynamic, workload-aware LLM serving framework based on morphological adaptation. MorphServe continuously monitors system load and morphs model components—transformer layers and KVC blocks—on the fly in response to real-time memory pressure. When resource usage surges, MorphServe reduces model footprint by replacing selected full-precision layers with lightweight quantized alternatives and expands KVC capacity by dynamically attaching additional memory blocks. These adaptations are reversed as pressure subsides, restoring full precision and reclaiming memory from KVC without interrupting inference.
@@
To achieve this, MorphServe introduces two complementary morphing mechanisms, both designed to support asynchronous and compatible kernel executions with minimal overhead: LayerSwapper identify low-impact transformer layers by a sensitivity-based profiling, selectively and asynchronously replacing them with lower-precision alternatives at runtime. KVResizer adaptively adjusts KVC capacity under memory pressure and runs in parallel with decoding using separate CUDA streams, ensuring seamless execution.
Across extensive experiments on Llama 2 [64], Llama 3 [22], CodeLlama [56], and Vicuna [63] using four datasets [27, 77, 23, 15] under Azure LLM Inference [4] and BurstGPT [67] traces, MorphServe reduces average SLO violations by 92.45% and P95 TTFT latency by 2.2$\times$–3.9$\times$ over full-precision serving, while preserving comparable accuracy. Compared to static quantization via AWQ [38], MorphServe reduces F1 and Rouge-L degradation by up to 88.85% and improves memory utilization by 29.29%. These results demonstrate MorphServe’s ability to adapt to dynamic workloads while balancing performance and responsiveness.
2 Background and Motivation
@@
Real-world LLM workloads are highly bursty. LLM serving systems face highly dynamic and bursty traffic patterns in real-world scenarios. As shown in Figure 1(a), the production workloads of Microsoft Azure LLM services [4, 61] and BurstGPT [67] reveal rapid fluctuations in both the request arrival rates (i.e., request bursts) and the volumes of tokens. These fluctuations reflect the non-stationary nature of practical LLM inference workloads, which deviates from the traditional assumptions of most serving schemes [33, 75, 30].
Request burst leads to long TTFT and SLO violation . As system load increases, even small surges can cause sharp spikes in time-to-first-token (TTFT) latency. In this work, we set the TTFT SLO threshold to 2 seconds, consistent with prior work [72, 20, 53]. As shown in Figure 1(b), full-precision serving quickly exceeds the SLO threshold once it reaches the saturation pointdefined as the load level at which available GPU memory becomes insufficient to schedule new requests for prefilling or to continue decoding for the ongoing batch. At this point, incoming requests are forced to wait until memory is reclaimed, incurring significant queueing latency with SLO violation.
Static quantization trades quality for efficiency irrespective of load. To mitigate resource constraints, static quantization methods [38, 17, 37] have been widely adopted. However, these methods introduce persistent accuracy degradation across all conditions, regardless of whether the system is overloaded. As shown in Figure 1(c), the INT4 quantized model with AWQ [38] consistently degrades accuracy—measured by F1 score following [37]—on the GovReport dataset from LongBench [5], even during periods when full-precision inference is feasible. This demonstrates that static quantization over-prioritizes efficiency, sacrificing model quality during low-load intervals.
Workload-aware adaptation achieves optimal tradeoffs. As shown in the Pareto analysis in Figure 1(d), MorphServe achieves superior tradeoffs by aligning dynamic quantization with real-time workload demand. A key insight is flexible, elastic mixed-precision LLM serving, where quantized and full-precision layers coexist and are dynamically reconfigured within a model in response to workload shifts. This contrasts with static quantization and recent dynamic methods [8, 18, 54], which prioritize serving performance or hardware efficiency but overlook runtime workload variability. Most importantly, MorphServe enables smooth navigation along the efficiency—accuracy Pareto frontier—from uncompressed, high-accuracy models to highly quantized, efficient ones.
@@ 3.1 Architecture and Workflow
System Architecture. As illustrated in Figure 2, MorphServe consists of three core components—Serving Monitor, Morphing Controller, and Morphing Actuator, which together form a feedback-driven control loop for dynamic adaptation.
@@
  • Morphing Controller serves as the global GPU memory manager. When monitored metrics exceed user-predefined thresholds (e.g., KVC memory usage > 85%, queueing delay > 100 ms), it decides whether to trigger selective layer swapping (Section 3.3) and elastic KVC resizing (Section 3.4), and dispatches corresponding instructions to the target workers.
  • Morphing Actuator resides on each worker and executes adaptation commands locally. It dynamically reconfigures the model using LayerSwapper (Section 3.3), which switches a selective set of layers between full-precision and pre-quantized layers, or between different quantization levels (e.g., from INT8 to INT4) to reduce resource usage and improve inference latency under pressure. In addition, it applies KVResizer (Section 3.4) to adjust KVC memory allocation by elastically expanding or shrinking the number of KVC blocks as needed. All adaptations are asynchronous overlapping communication and computation [60] with preallocated memory buffers to seamlessly overlap with ongoing inference.
@@
Token-level Workload Adaptation. Unlike existing model and KVC compression schemes, which affect the entire request [52, 38, 17, 37, 76], MorphServe enables fine-grained, token-level workload adaptation. During a single request’s decoding phase, MorphServe may temporarily replace a subset of layers. For example, switching 2 layers from full-precision to INT4 when saturation is detected (examples shown in Section 3.3). This allows early tokens to be generated at full precision, while only later tokens experience minimal accuracy degradation. Once the pressure subsides, the affected layers are restored to full precision, enabling continued decoding at the original accuracy. As a result, accuracy degradation is confined to a small portion of tokens, even within a single request.
@@ 3.2 Offline Profiling for Layer Swapping Sequence
Let $f(x_{t})$ denote the full-precision model output at time $t$, and $f^{(Q_{t})}(x_{t})$ the output when a subset of layers $Q_{t}$ are quantized at that time. The cumulative degradation over the interval $[t_{1},t_{n}]$ can be formulated as:
@@
In this formulation, $\mathrm{LTS}_{p}$ and $\mathrm{LRS}_{p}$ are local sensitivity metrics that evaluate the behavior of the layer $p$ in isolation, while $\mathrm{MDS}^{(Q)}_{p}$ is a global metric that measures the model-level degradation when replacing $p$, given the current replaced layer set $\mathcal{Q}$. For a given model, the LIS for each layer is computed offline during profiling, and the resulting sequence is stored and used directly at runtime. This design avoids any runtime recomputation or decision-making overhead. Full details on the scoring, hyperparameter tuning, and selection algorithms are provided in the Appendix.
@@ 3.3 LayerSwapper: Runtime Layer Swapping
  • Model Preloading and Kernel Precompilation. Prior to serving, all decoder layer variants (e.g., FP16, INT8, INT4, and INT3) are preloaded into a contiguous, pinned CPU memory region, while the full-precision model replica is loaded into a preallocated contiguous GPU memory, as shown in the Figure 2. MorphServe tracks the memory addresses of all layer variants, enabling efficient direct memory copies for layer swapping. To avoid runtime latency, inference kernels corresponding to precision levels are precompiled in advance. We also implement kernel fusion to optimize performance, while the rest of the serving pipeline reuses state-of-the-art techniques—such as PagedAttention [33] and FlashAttention [12, 11]—to ensure compatibility and efficiency.
  • Asynchronous In-place Layer Swapping. At runtime, MorphServe performs in-place layer swapping using asynchronous CUDA streams to avoid interference with ongoing decoding. As illustrated in Figure 3, when layers 28 and 29 are selected for replacement, the swapping process is launched asynchronously while earlier layers (e.g., 0–27) continue computation without interruption. Full-precision layers are safely discarded from GPU memory since their backup copies reside in pinned CPU memory, and quantized variants are copied into the same memory addresses to avoid pointer remapping. Due to the relatively compact size of each decoder layer (e.g.,  0.4 GB for FP16 and  0.1 GB for INT4 in Llama 2 7B), the PCIe transfer latency is minimal - approximately 4 ms for INT4 and 16 ms for FP16 for Llama2 7B on PCIe Gen4 with up to 26-28 GB/s bandwidth. In practice, the complete layer swapping process for a INT4 variant—including memory transfer and reconstruction—takes approximately 6 ms and is fully overlapped with decoding, resulting in negligible TPOT overhead. Additional performance breakdowns are provided in Section 4.

### v2

@@ Abstract
Efficiently serving large language models (LLMs) under dynamic and bursty workloads remains a key challenge for real-world deployment. Existing serving frameworks and static model compression techniques fail to adapt to workload fluctuations, leading to either service-level objective (SLO) violations under full-precision serving or persistent accuracy degradation with static quantization. To deal with these issues, we present MorphServe, a dynamic, workload-aware LLM serving framework based on morphological adaptation. MorphServe introduces two asynchronous, token-level runtime mechanisms: quantized layer swapping, which selectively replaces less impactful layers with quantized alternatives during high-load periods, and pressure-aware KV cache resizing, which dynamically adjusts KV cache capacity in response to memory pressure. These mechanisms enable state-preserving transitions with minimum runtime overhead and are fully compatible with modern scheduling and attention techniques. Extensive experiments on Vicuna and Llama family models with real-world workloads demonstrate that MorphServe reduces average SLO violations by 92.45% and improves the P95 TTFT latency by $2.2$-$3.9\times$ compared to full-precision serving, without compromising generation quality. These results establish MorphServe as a practical and elastic solution for LLM deployment in dynamic environments.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
The rise of large language models (LLMs) has made efficient and reliable serving a core challenge in modern AI infrastructure. Systems like vLLM Kwon et al. (2023) and Orca Yu et al. (2022) optimize throughput via PagedAttention Kwon et al. (2023) and continuous batching Yu et al. (2022); Sun et al. (2024); He et al. (2024b), but assume fixed-precision execution and stable workloads. In contrast, real-world LLM workloads are dynamic and bursty Wang et al. (2024); Patel et al. (2024); Azure (2024), with fluctuating request rates and context lengths. Even brief load spikes can cause memory exhaustion or queueing delays, leading to SLO violations—such as increased time-to-first-token (TTFT) and time-per-output-token (TOPT)—that degrade both user experience and system throughput.
@@
One naive solution is to statically over-provision GPU resources to accommodate worst-case traffic spikes. However, over-provisioning leads to substantial cost inefficiencies during underutilized periods Jaiswal et al. (2025); Fu et al. (2024). Moreover, edge deployments lack the flexibility for dynamic scaling altogether Cai et al. (2024a). Thus, the inability to elastically match model resource usage to real-time demand results in either SLO violations under pressure, or significant resource waste during low-load intervals.
@@
Model compression techniques, such as quantization Lin et al. (2024b); Frantar et al. (2022); Lin et al. (2024a); Shao et al. (2023), pruning Ma et al. (2023); Sreenivas et al. (2024); Gao et al. (2024b), or low-rank approximation Hu et al. (2022); Wu et al. (2024), offer an alternative approach by statically reducing the resource footprint of deployed LLMs. While these methods are effective in lowering memory and compute demands, they introduce irreversible accuracy degradation that persists even during periods of low load, when full-precision inference could be served without penalty. This results in a rigid, suboptimal quality–efficiency tradeoff that fails to align with workload variability. Key-value cache (KVC) compression Cai et al. (2024b); Zhou et al. (2024); Li et al. (2024b) and eviction Liu et al. (2023); Feng et al. (2024) methods have been proposed to further reduce memory usage. However, these techniques often rely on fixed heuristics, cannot adapt to different workloads, lack compatibility with modern attention variants like Grouped Query Attention (GQA) Ainslie et al. (2023); Chen et al. (2024b); Chinnakonduru and Mohapatra (2024) and Multi-Head Latent Attention (MLA) Meng et al. (2025); Liu et al. (2024), and remain inflexible to runtime serving conditions.
@@
To handle dynamic workloads and address the above issues, in this paper, we present MorphServe, a dynamic, workload-aware LLM serving framework based on morphological adaptation. MorphServe continuously monitors system load and morphs model components—transformer layers and KVC blocks—on the fly in response to real-time memory pressure. When resource usage surges, MorphServe reduces model footprint by replacing selected full-precision layers with lightweight quantized alternatives and expands KVC capacity by dynamically attaching additional memory blocks. These adaptations are reversed as pressure subsides, restoring full precision and reclaiming memory from KVC without interrupting inference.
@@
To achieve this, MorphServe introduces two complementary morphing mechanisms, both designed to support asynchronous and compatible kernel executions with minimal overhead: LayerSwapper identifies low-impact transformer layers by a sensitivity-based profiling, selectively and asynchronously replacing them with lower-precision alternatives at runtime. KVResizer adaptively adjusts KVC capacity under memory pressure and runs in parallel with decoding using separate CUDA streams, ensuring seamless execution.
[…]
2 Related Work
LLM Serving Systems. TorchServe PyTorch (2023) and NVIDIA Triton NVIDIA Corporation (2019b) provide general-purpose serving frameworks for parallel inference. Recent systems specialize in LLM serving NVIDIA Corporation (2019a); Hugging Face , improving scheduling and memory management. Orca Yu et al. (2022) introduces continuous batching, and vLLM Kwon et al. (2023) proposes PagedAttention, both aimed at improving throughput. Sarathi Agrawal et al. (2024) and FastServe Wu et al. (2023) further reduce queuing delays via chunked prefill or preemptive scheduling. Despite these advances, existing systems assume fixed precision and static memory provisioning, which limits their responsiveness to bursty workloads. In contrast, MorphServe enables fully online, workload-aware adaptation through runtime quantized layer swapping and elastic KV cache resizing, achieving smooth performance-accuracy tradeoffs under overload.
LLM Post-Training Quantization. Post-training quantization (PTQ) reduces inference cost by compressing model weights Dettmers et al. (2023); Kim et al. (2023); Li et al. (2024a); Zhao et al. (2024); Lin et al. (2024b); Frantar et al. (2022); Xiao et al. (2023); Lin et al. (2024a). Round-to-nearest schemes Yao et al. (2023); Nagel et al. (2021) are simple but inaccurate for outlier-heavy layers. Non-uniform and calibration-based methods, including AdaRound Nagel et al. (2020), ZeroQuant Yao et al. (2022), GPTQ Frantar et al. (2022), and AWQ Lin et al. (2024b), achieve higher fidelity via adaptive rounding or activation statistics. Activation quantization has also been explored (e.g., SmoothQuant Xiao et al. (2023), QServe Lin et al. (2024c), DuQuant Lin et al. (2024a)), but remains static and incurs irreversible degradation even when resources permit full precision. MorphServe provides a systematic solution that dynamically switches between different precision layers based on real-time service pressure, compatible with any weight quantization method, such as GPTQ and AWQ.
Mixed-Precision and Adaptive-Depth Inference. Mixed-precision execution improves efficiency by leveraging lower-bit operations Reggiani et al. (2023); Huang et al. (2021b); Wang et al. (2020); Chen et al. (2024a); Frantar et al. (2025); Dong et al. (2019); Risso et al. (2022). Hardware-level methods (Mix-GEMM Reggiani et al. (2023), Bit-Split Wang et al. (2020)) and prompt-aware schemes (PMPD Chen et al. (2024a), MARLIN Frantar et al. (2025)) adjust precision based on token entropy or pre-computed sensitivity. Recent adaptive-depth and early-exit frameworks (e.g., FlexiDepth Luo et al. (2025), LayerSkip Elhoushi et al. (2024)) reduce computation via layer skipping but require architectural modification and lack real-time adaptation. In contrast, MorphServe achieves dynamic, workload-aware mixed-precision serving guided by runtime telemetry, coupling token-level precision with system load for smooth efficiency–accuracy balancing.
3 Motivation
@@
Real-world LLM workloads are highly bursty. LLM serving systems face highly dynamic and bursty traffic patterns in real-world scenarios. As shown in Figure 1(a), the production workloads of Microsoft Azure LLM services Azure (2024); Stojkovic et al. (2025) and BurstGPT Wang et al. (2024) reveal rapid fluctuations in both the request arrival rates (i.e., request bursts) and the volumes of tokens. These fluctuations reflect the non-stationary nature of practical LLM inference workloads, which deviates from the traditional assumptions of most serving schemes Kwon et al. (2023); Yu et al. (2022); Jaiswal et al. (2025).
Request burst leads to long TTFT and SLO violation. As system load increases, even small surges can cause sharp spikes in TTFT latency. In this work, we set the TTFT SLO threshold to 2 seconds, consistent with prior work Xiong et al. (2024); Gao et al. (2024a); Qiao et al. (2024). As shown in Figure 1(b), full-precision serving quickly exceeds the SLO threshold once it reaches the saturation point: defined as the load level at which GPU memory becomes insufficient to schedule new requests for prefilling or to continue decoding for the ongoing batch. At this point, incoming requests are forced to wait until memory is reclaimed, incurring significant queueing latency with SLO violation.
Static quantization trades quality for efficiency irrespective of load. To mitigate resource constraints, static quantization methods Lin et al. (2024b); Frantar et al. (2022); Lin et al. (2024a) have been widely adopted. However, these methods introduce persistent accuracy degradation across all conditions, regardless of whether the system is overloaded. As shown in Figure 1(c), the INT4 quantized model with AWQ Lin et al. (2024b) consistently degrades accuracy, measured by F1 score following Lin et al. (2024a), on the GovReport dataset from LongBench Bai et al. (2023), even during periods when full-precision inference is feasible. This demonstrates that static quantization over-prioritizes efficiency, sacrificing quality during low-load intervals.
Workload-aware adaptation achieves optimal tradeoffs. As shown in the Pareto analysis in Figure 1(d), MorphServe achieves superior tradeoffs by aligning dynamic quantization with real-time workload conditions. A key insight in MorphServe is adaptive mixed-precision LLM serving, where quantized and full-precision layers can coexist and are dynamically reconfigured. This contrasts with static quantization and recent dynamic methods Chen et al. (2024a); Frantar et al. (2025); Reggiani et al. (2023), which prioritize serving performance or hardware efficiency but overlook runtime workload variability. Most importantly, MorphServe enables smooth navigation along the accuracy—efficiency Pareto frontier, from uncompressed, high-accuracy models to highly quantized, efficient ones.
@@ 4.1 Architecture and Workflow
System Architecture. As illustrated in Figure 2, MorphServe consists of three core components—Serving Monitor, Morphing Controller, and Morphing Executor, which together form a feedback-driven control loop for dynamic adaptation.
@@
  • Morphing Controller serves as the global GPU memory manager. When monitored metrics exceed user-predefined thresholds (e.g., KVC memory usage over 85%, queueing delay over 100 ms), it decides whether to trigger selective layer swapping (Section 4.3) and elastic KVC resizing (Section 4.4), and dispatches corresponding instructions to the target workers.
  • Morphing Executor resides on each worker and executes adaptation commands locally. It dynamically reconfigures the model using LayerSwapper (Section 4.3), which switches a selective set of layers between full-precision and pre-quantized layers, or between different quantization levels (e.g., from INT8 to INT4) to reduce resource usage and improve inference latency under pressure. In addition, it applies KVResizer (Section 4.4) to adjust KVC memory allocation by elastically expanding or shrinking the number of KVC blocks as needed. All adaptations are asynchronous overlapping communication and computation Steve Rennich with preallocated memory buffers to seamlessly overlap with ongoing inference.
@@
Token-level Workload Adaptation. Unlike existing model and KVC compression schemes, which affect the entire request Polino et al. (2018); Lin et al. (2024b); Frantar et al. (2022); Lin et al. (2024a); Zhao et al. (2024), MorphServe enables fine-grained, token-level workload adaptation. During a single request’s decoding phase, MorphServe may temporarily replace a subset of layers. For example, switching 2 layers from full-precision to INT4 when saturation is detected (examples shown in Section 4.3). This allows early tokens to be generated at full precision, while only later tokens experience minimal accuracy degradation. Once the pressure subsides, the affected layers are restored to full precision, enabling continued decoding at the original accuracy. As a result, accuracy degradation is confined to a small portion of tokens, even within a single request.
@@ 4.2 Offline Profiling for Layer Swapping Sequence
Let $f(x_{t})$ denote the full-precision model output at time $t$, and $f^{(Q_{t})}(x_{t})$ the output when a subset of layers $Q_{t}$ are quantized at that time. The cumulative degradation over the interval $t_{1},t_{n}]$ can be formulated as:
@@
Independence of Layer Quantization Effects. To validate the independence of layer-wise quantization effects, we incrementally quantized the early layers of Llama 2 7B. We observed that the additional perplexity from quantizing a layer remained nearly constant across all settings. This supports the assumption that quantization impact is approximately additive across layers. For detailed results, please refer to Appendix Table 4.
@@
In this formulation, $\mathrm{LTS}_{p}$ and $\mathrm{LRS}_{p}$ are local sensitivity metrics that evaluate the behavior of the layer $p$ in isolation, while $\mathrm{MDS}^{(Q)}_{p}$ is a global metric that measures the model-level degradation when replacing $p$, given the current replaced layer set $\mathcal{Q}$. For a given model, the LIS for each layer is computed offline during profiling, and the resulting sequence is stored and used directly at runtime. This design avoids any runtime recomputation or decision-making overhead. We use a weighted combination of three sensitivity metrics to compute the LIS: weight sensitivity ($\alpha_{1}$), quantization distortion ($\alpha_{2}$), and end-to-end degradation ($\beta$), with $\alpha_{1}=\alpha_{2}=0.25$ and $\beta=0.5$. This configuration emphasizes stability and user-perceived quality while retaining fine-grained layer-local signals. Compared to a uniform weighting scheme, our setting yields more consistent layer rankings and improved performance. Full results and ablations appear in Appendix Table 1.
Metric Choice in Sensitivity Computation. We use cosine similarity instead of the L2 norm to assess layer sensitivity, as it captures direction-based shifts across transformer layers and mitigates scale noise. This choice follows analyses of layer relevance in Transformers He et al. (2024a); Sun et al. (2025); Jiang et al. (2025). Cosine-based LIS yields lower perplexity than L2 variants across quantization levels, confirming its suitability for layer profiling. Our detailed evaluation comparison is in Appendix Table 2.
@@ 4.3 LayerSwapper: Runtime Layer Swapping
  • Model Preloading and Kernel Precompilation. Prior to serving, all decoder layer variants (e.g., FP16, INT8, INT4, and INT3)111“INT#” denotes weight-only quantization (W#). are preloaded into a contiguous, pinned CPU memory region, while the full-precision model replica is loaded into a preallocated contiguous GPU memory, as shown in the Figure 2. MorphServe tracks the memory addresses of all layer variants, enabling efficient direct memory copies for layer swapping. To avoid runtime latency, inference kernels corresponding to precision levels are precompiled in advance. We also implement kernel fusion to optimize performance, while the rest of the serving pipeline reuses state-of-the-art techniques—such as PagedAttention Kwon et al. (2023) and FlashAttention Dao et al. (2022); Dao (2023)—to ensure compatibility and efficiency.
  • Asynchronous In-place Layer Swapping. At runtime, MorphServe performs in-place layer swapping using asynchronous CUDA streams to avoid interference with ongoing decoding. As illustrated in Figure 3, when layers 28 and 29 are selected for replacement, the swapping process is launched asynchronously while earlier layers (e.g., 0–27) continue computation without interruption. Full-precision layers are safely discarded from GPU memory since their backup copies reside in pinned CPU memory, and quantized variants are copied into the same memory addresses to avoid pointer remapping. Due to the relatively compact size of each decoder layer (e.g.,  0.4 GB for FP16 and  0.1 GB for INT4 in Llama 2 7B), the PCIe transfer latency is minimal - approximately 4 ms for INT4 and 16 ms for FP16 for Llama2 7B on PCIe Gen4 with up to 26-28 GB/s bandwidth. In practice, the complete layer swapping process for an INT4 variant—including memory transfer and reconstruction—takes approximately 6 ms and is fully overlapped with decoding, resulting in negligible TPOT overhead. Additional performance breakdowns are provided in Section 5.
Figure 4: MorphServe provides the best latency–accuracy tradeoff across four models and two traces, with four datasets. MorphServe in accuracy mode (dark green) reduces P95 TTFT by $2.2\times$$3.9\times$ compared to full-precision serving while maintaining comparable generation quality. In performance mode (light green), MorphServe consistently outperforms INT4 quantized models in output quality with no additional latency overhead.